Giving a ball tide to a string, length 8 cm. And rotated in a circular motion around a central axis, with an Omga of Pi over 4 radiant with a constant speed of 4 meters per second- in what pattern will the ball fly- off, if the string will be cut off when it's perpendicular to the horizon?


1
Expert's answer
2021-03-31T09:31:57-0400
"x=v_xt=4t\\y=h-0.5gt^2=h-4.9t^2"